# Optical-Flow Guided Prompt Optimization for Coherent Video Generation

### v1

@@ 1 Introduction
Recently, diffusion models have become the de-facto standard for image generation. In particular, text-to-image (T2I) diffusion models [dhariwal2021diffusion, rombach2022high] have gained significant attention due to their ability to enable users control over the generation process via text prompts.
Building on these advancements, research has now progressed toward text-to-video (T2V) diffusion models [chen2024videocrafter2, guo2023animatediff, wang2023lavie] that generate videos which are both visually engaging and contextually coherent. Despite these advances, achieving temporally consistent videos remains a challenge in many T2V models. Although recent work has attempted to address this, many proposed methods rely on fine-tuning [ge2023preserve] or introduce unnecessary interventions that can potentially adversely impact [chen2024unictrl, lee2024videoguide].
@@ 2 Related Works
Video Latent Diffusion Models. Video latent diffusion models [he2022latent] have gained attraction for their ability to efficiently generate videos by operating within a compressed latent space, thereby reducing the computational cost and memory demands associated with high-resolution video generation. Building on this approach, VideoCrafter [chen2023videocrafter1, chen2024videocrafter2], AnimatedDiff [guo2023animatediff] and Lavie [wang2023lavie] extended latent diffusion to handle text-to-video generation to produce contextually relevant and controllable video outputs. While these advancements have significantly improved video generation, ensuring temporal consistency remains challenging. FreeInit [wu2023freeinit] alleviates this problem by refining the initial noise to ensure low-frequency information. While temporal consistency is increased, the process of iteratively obtaining clean videos is computationally expensive and can result in a loss of detail. UniCtrl [chen2024unictrl] and VideoGuide [lee2024videoguide] address this issue by introducing attention injection and leveraging different pre-trained video diffusion models, respectively. However, these approaches can sometimes negatively impact performance: UniCtrl’s injection mechanism struggles with handle color changes, while VideoGuide depends on the performance of the additional pre-trained model. Nevertheless, these methods are orthogonal to ours, as we optimize the prompt, and can be combined to further enhance video generation quality.
@@ 3 Main Contribution: MotionPrompt
3.1 Conditional Video Diffusion
Video latent diffusion models encode $N$-frame clean video, $\{{\bm{x}}^{(i)}\}_{i=1}^{N}\in{\mathbb{R}}^{N\times C\times H\times W}$, to $\{{\bm{z}}_{0}^{(i)}\}_{i=1}^{N}$ using an encoder $\mathcal{E}$, where $C,H$ and $W$ represent channel, height and width of the video, respectively. Unless otherwise noted, we simply denote ${\bm{z}}_{0}$ as $\{{\bm{z}}_{0}^{(i)}\}_{i=1}^{N}$ for convenience and ${\bm{z}}_{0}\sim p_{0}({\bm{z}})$.
@@ 3.1 Conditional Video Diffusion
${\bm{z}}_{t-1}$ $=\sqrt{\alpha_{t-1}}\hat{{\bm{z}}}_{t}+\sqrt{1-\alpha_{t-1}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (3)
$\hat{{\bm{z}}}_{t}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})\right),$ (4)
@@ 3.2 Prompt Optimization for Video Guidance
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{t}(\hat{\bm{c}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\hat{\bm{c}}_{t})$ (9)
$\hat{{\bm{z}}}_{t}(\hat{\bm{c}}_{t})$ $=\left({\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\hat{\bm{c}}_{t})\right)/{\sqrt{\alpha_{t}}}$ (10)
@@ 3.3 Defining Loss Function from Optical Flow
TV loss for optical flow. Additionally, to ensure that the optical flow adheres to the assumption of a smooth field, we incorporate total variation (TV) loss as an additional regularization term. For simplicity, in the following equation, we will denote ${\mathbf{f}}(\hat{{\bm{x}}}_{t}({\bm{c}}({\mathcal{T}}))$ as $\mathbf{f}$. This TV loss is then defined as
@@ 4.1 Implementation Details
Discriminator training. For discriminator training, we sample videos from each model using the same set of 800 prompts. Leveraging our prompt optimization, the discriminator operates on single images rather than full videos. We fine-tune a pre-trained Vision Transformer (ViT) [dosovitskiy2021an] as an image encoder, adding a projection layer to adapt two-channel optical flow for ViT and using a 3-layer MLP as the classifier. This setup achieved rapid convergence in under 20 epochs. Optical flow was extracted with RAFT [teed2020raft] from real videos selected from the DAVIS [pont20172017] and WebVid [bain2021frozen], as well as from generated videos. Finally, we train and deploy a separate discriminator for each video model. All training and experiments were conducted on a single 40GB NVIDIA A100 GPU.
@@ 4.2 Results
For quantitative comparison, we evaluate five key metrics from VBench [huang2023vbench]—subject consistency, background consistency, temporal flickering, motion smoothness, and dynamic degree—to assess improvements in consistency and motion. We also measure overall consistency to confirm that prompt optimization maintains fidelity to the text prompt. Table 1 shows that our method improves object consistency, reduces temporal flickering, and enhances motion smoothness with minimal impact on text alignment. We are aware that there may exist a trade-off between consistency and motion dynamics in the proposed method. However, visualization results demonstrate that our method balances dynamics and coherence effectively. See supplementary materials for more details.
@@
Additionally, we conduct a user study to evaluate our method’s effectiveness in overall quality, temporal quality, and text alignment. We select three videos from each of the three models, totaling nine videos, and 20 participants rated them on a scale of 1 to 5. As shown in Table 2, our method achieved higher ratings in all aspects, demonstrating its ability to produce visually appealing, temporally consistent, and well-aligned videos.
Aspects Baselines Ours
Overall Quality 2.65 4.32
Temporal Quality 2.71 4.29
Text Alignment 3.33 4.56
Table 2: User study result. Bold: Best
@@ 5 Additional Experiments
We performed additional analysis to support the effectiveness of our method and to analyze the design components of the proposed approach, focusing on the AnimateDiff [wang2023lavie].
5.1 Token Variations
@@ 5.2 Ablation Study
In Table 3, we examine how key hyperparameters affect our framework’s performance, focusing on the optimal number of iterations per optimization step and the most effective stage within the 50-step sampling process. For iteration count, a single iteration was insufficient, while 7 iterations smoothed motion but reduced dynamic degree and overall consistency, as well as increasing computation time. An effective balance was achieved with 3 iterations. Similarly, starting optimization early improved motion smoothness but reduced consistency, while delaying it too late weakened its impact. Based on these findings, we established an optimal range for applying optimization. Except for the hyperparameters being compared, all other settings use the optimal hyperparameters listed in the supplementary materials.
@@
Analysis of each loss component. We analyze our loss components to demonstrate the effectiveness of each in Fig. 5. We observed that $\ell_{disc}$ effectively promotes natural motion and regulates excessively changing objects. However, some objects still fail to maintain consistency, or they remain as residual artifacts. When we additionally apply $\ell_{TV}$, i.e., our final loss $\ell_{total}$ , these issues are significantly reduced, allowing us to reliably generate cleaner videos.
Figure 5: Analysis of each loss component. Our final loss demonstrates improved appearance fidelity and temporal coherence of motion relative to the original video.
@@ A.2 Hyperparameters for Evaluation
Table 4 lists the hyperparameters used for our quantitative evaluation. The ‘# of frames’ denotes the number of video frames decoded into pixels and used for optical flow calculation. Specifically, if the value is 6, this indicates that 3 sets of adjacent frames were sampled, resulting in 3 optical flows being used to compute the corresponding loss.
@@ A.3 User study
For the user study, we utilize videos presented in the paper and on the project webpage. Participants are asked to evaluate the videos based on the following questions: (1) Overall Quality: Does the generated video exhibit good overall quality? (2) Motion Smoothness: Are the motions and transitions in the generated video smooth? (3) Text Alignment: Does the generated video align well with the given textual conditions?
@@ B.1 Ablation on Token Optimization
Here, we conduct an ablation study to investigate the impact of various configurations related to token optimization on the overall performance (Table 5). Except for the factors being examined, all other settings were kept consistent with the values in Table 4.
@@
B.2 Exploring Discriminator Generalization
We primarily use a discriminator trained on paired data. In this setup, the model generating fake data for discriminator training matches the model used during inference. To explore the impact of the discriminator’s training approach and evaluate its generalization capability, we conduct inference using a discriminator trained on a different dataset, referred to as cross-dataset inference.
Table 6 presents the results of this evaluation. Surprisingly, we observe a general improvement in performance when using a discriminator trained on different data (i.e., videos generated by Lavie [wang2023lavie] and VideoCrafter2 [chen2024videocrafter2]). We conjecture that this is because the baseline performance of Lavie and VideoCrafter2 is higher compared to AnimateDiff, making it more challenging for the discriminator to differentiate these videos from real ones. This likely resulted in stricter training, which may have contributed to the improved performance observed during inference. These findings suggest the potential for further performance enhancements through improved discriminator training strategies.
Source Model for Fake Data AD (Defualt) Lavie VC2
Subject Consistency 0.9528 0.9625 0.9535
Background Consistency 0.9763 0.9753 0.9764
Temporal Flickering 0.9258 0.9490 0.9283
Motion Smoothness 0.9599 0.9691 0.9617
Dynamic Degree 0.4125 0.4088 0.4100
Overall Consistency 0.2529 0.2473 0.2509
B.3 Synergies with the Existing Method
@@
While FreeInit significantly improves temporal quality, it does so at the expense of overall video quality. Specifically, Table 7 demonstrates that increasing the number of initialization steps leads to a sharp decline in metrics related to video generation quality, such as Overall Consistency and Image Quality. Fig. 7 further reveal that, compared to the baseline, the videos generated with FreeInit often lose high-frequency details and exhibit noticeable saturation. However, by applying our method after a single noise initialization step, we observed an improvement in temporal quality while relatively minimizing the compromise in video quality, compared to the FreeInit approach where noise initialization is applied multiple times.

### v2

@@ 1 Introduction
Recently, diffusion models have become the de-facto standard for image generation. In particular, text-to-image (T2I) diffusion models [dhariwal2021diffusion, rombach2022high] have gained significant attention due to their ability to enable users control over the generation process via text prompts. Building on these advancements, research has now progressed toward text-to-video (T2V) diffusion models [chen2024videocrafter2, guo2023animatediff, wang2023lavie] that generate videos which are both visually engaging and contextually coherent. Despite these advances, achieving temporally consistent videos remains a challenge in many T2V models. Although recent work has attempted to address this, many proposed methods rely on fine-tuning [ge2023preserve] or introduce unnecessary interventions that can potentially adversely impact [chen2024unictrl, lee2024videoguide].
@@ 2 Related Works
Latent Video Diffusion Models. Latent video diffusion models [he2022latent, hong2023cogvideo, yang2024cogvideox, blattmann2023align] have gained attraction for their ability to efficiently generate videos by operating within a compressed latent space, thereby reducing the computational cost and memory demands associated with high-resolution video generation. Building on this approach, VideoCrafter [chen2023videocrafter1, chen2024videocrafter2], AnimatedDiff [guo2023animatediff] and Lavie [wang2023lavie] extended latent diffusion to handle text-to-video generation to produce contextually relevant and controllable video outputs. While these advancements have significantly improved video generation, ensuring temporal consistency remains challenging. FreeInit [wu2023freeinit] alleviates this problem by refining the initial noise to ensure low-frequency information. While temporal consistency is increased, the process of iteratively obtaining clean videos is computationally expensive and can result in a loss of detail. UniCtrl [chen2024unictrl] and VideoGuide [lee2024videoguide] address this issue by introducing attention injection and leveraging different pre-trained video diffusion models, respectively. However, these approaches can sometimes negatively impact performance: UniCtrl’s injection mechanism struggles with handle color changes, while VideoGuide depends on the performance of the additional pre-trained model. Nevertheless, these methods are orthogonal to ours, as we optimize the prompt, and can be combined to further enhance video generation quality.
@@ 3 Main Contribution: MotionPrompt
3.1 Conditional Video Diffusion Model
Latent video diffusion models encode $N$-frame clean video, $\{{\bm{x}}^{(i)}\}_{i=1}^{N}\in{\mathbb{R}}^{N\times C\times H\times W}$, to $\{{\bm{z}}_{0}^{(i)}\}_{i=1}^{N}$ using an encoder $\mathcal{E}$, where $C,H$ and $W$ represent channel, height and width of the video, respectively. Unless otherwise noted, we simply denote ${\bm{z}}_{0}$ as $\{{\bm{z}}_{0}^{(i)}\}_{i=1}^{N}$ for convenience and ${\bm{z}}_{0}\sim p_{0}({\bm{z}})$.
@@ 3.1 Conditional Video Diffusion Model
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{z}}}_{t}+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (3)
$\hat{{\bm{z}}}_{t}$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left({\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})\right),$ (4)
@@ 3.2 Prompt Optimization for Video Guidance
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{z}}}_{t}(\hat{\bm{c}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\hat{\bm{c}}_{t})$ (9)
$\hat{{\bm{z}}}_{t}(\hat{\bm{c}}_{t})$ $=\left({\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\hat{\bm{c}}_{t})\right)/{\sqrt{\bar{\alpha}_{t}}}$ (10)
@@ 3.3 Defining Loss Function from Optical Flow
TV loss for optical flow. Additionally, to ensure that the optical flow adheres to the assumption of a smooth field, we incorporate total variation (TV) loss as an additional regularization term. For simplicity, in the following equation, we will denote ${\mathbf{f}}(\hat{{\bm{x}}}_{t}({\bm{c}}({\mathcal{T}})))$ as $\mathbf{f}$. This TV loss is then defined as
@@ 4.1 Implementation Details
Discriminator training. For discriminator training, we sample videos from each model using the same set of 800 prompts. We highlight that the discriminator is trained on the optical flow of complete, clean videos, whereas during inference, it processes the optical flow of denoised estimates. This distinction ensures that, even with the same prompts, the discriminator encounters unseen data. We fine-tune a pre-trained Vision Transformer (ViT) [dosovitskiy2021an] as an image encoder, adding a projection layer to adapt two-channel optical flow for ViT and using a 3-layer MLP as the classifier. This setup achieved rapid convergence in under 20 epochs. Optical flow was extracted with RAFT [teed2020raft] from real videos selected from the DAVIS [pont20172017] and WebVid [bain2021frozen], as well as from generated videos. Finally, we train and deploy a separate discriminator for each video model. All training and experiments were conducted on a single 40GB NVIDIA A100 GPU.
@@ 4.2 Results
For quantitative comparison, we evaluate five key metrics from VBench [huang2023vbench]—subject consistency, background consistency, temporal flickering, motion smoothness, and dynamic degree—to assess improvements in consistency and motion. We also measure overall consistency to confirm that prompt optimization maintains fidelity to the text prompt. Tab. 1 shows that our method improves object consistency, reduces temporal flickering, and enhances motion smoothness with minimal impact on text alignment. We are aware that there may exist a trade-off between consistency and motion dynamics in the proposed method. However, visualization results demonstrate that our method balances dynamics and coherence effectively. See supplementary materials for more details.
@@
Additionally, we conduct an A/B user study. We showed videos generated by the baseline and our method for the same prompts, asking participants to choose the one with better temporal quality.
Baseline Win Tie Lose
AnimateDiff 66.5 17.8 15.7
Lavie 55.1 21.1 23.8
VideoCrafter2 53.0 17.7 29.3
Table 2: User study results.
Among 100 participants evaluating a total of 90 videos (30 videos per model), our method was preferred overall (Tab. 2).
@@ 5 Additional Experiments
We performed additional analysis to analyze the design components of the proposed approach and to support the effectiveness of our method, focusing on the AnimateDiff [wang2023lavie].
@@ 5.1 Ablation Study
Analysis of each loss component.
In  Tab. 3, we examine how key hyperparameters affect our framework’s performance. First, we validate the effectiveness and robustness of our loss components. ($\lambda_{1}$) A comparison between rows (a) and (c) reveals that incorporating $\ell_{disc}$ further improves the results. ($\lambda_{2}$) To investigate the role of $\lambda_{2}$, we gradually increased its value. As shown in rows (b)–(d), a higher weight results in smoother motion but comes at the cost of performance in other metrics. Based on this observation, we selected a moderate weight to balance the trade-off. ($\lambda_{3}$) This loss was introduced after identifying a degradation in video quality, which we hypothesized was due to the optimized embedding deviating from the text space the diffusion model was trained on. Comparing rows (c) and (e) supports this hypothesis, as certain metrics show a decline. While performance varies with different hyperparameter choices, all configurations consistently outperform the baseline, demonstrating that our method remains robust and is not overly sensitive to hyperparameter selection.
@@
We also explored the optimal number of iterations per optimization step and the most effective stage within the 50-step sampling process. For iteration count, a single iteration was insufficient, while 7 iterations smoothed motion but reduced dynamic degree and overall consistency, as well as increasing computation time. An effective balance was achieved with 3 iterations. Similarly, starting optimization early improved motion smoothness but reduced consistency, while delaying it too late weakened its impact. Based on these findings, we established an optimal range for applying optimization. Except for the hyperparameters being compared, all other settings use the optimal hyperparameters listed in the supplementary materials.
Loss Weight
$\lambda_{1}$ / $\lambda_{2}$ / $\lambda_{3}$ Subject Consistency ($\uparrow$) Motion Smoothness ($\uparrow$) Dynamic Degree ($\uparrow$) Overall Consistency ($\uparrow$)
(a) 0.0 / 5.0 / 10.0 0.9504 0.9643 0.4365 0.2528
(b) 1.0 / 0.0 / 10.0 0.9522 0.9648 0.4232 0.2522
(c) 1.0 / 5.0 / 10.0 0.9528 0.9599 0.4125 0.2529
(d) 1.0 / 10.0 / 10.0 0.9514 0.9658 0.4072 0.2525
(e) 1.0 / 5.0 / 0.0 0.9505 0.9650 0.4232 0.2517
@@
5.2 Discriminator Generalization
We primarily use a discriminator trained on paired data, where the model generating fake data matches the one used during inference. To assess the discriminator’s training impact and generalization, we also perform cross-dataset inference using a discriminator trained on a different dataset. Tab. 4 presents the results of this evaluation. Surprisingly, we observe a general improvement in performance when using a discriminator trained on different data (i.e., videos generated by Lavie and VideoCrafter2). We conjecture that this is because the baseline performance of Lavie and VideoCrafter2 is higher compared to AnimateDiff, making it more challenging for the discriminator to differentiate these videos from real ones. This likely resulted in stricter training, which may have contributed to the improved performance observed during inference. These findings suggest the potential for further performance enhancements through improved discriminator training strategies.
Source Model for Fake Data AD (Defualt) Lavie VC2
Subject Consistency 0.9528 0.9625 0.9535
Background Consistency 0.9763 0.9753 0.9764
Temporal Flickering 0.9258 0.9490 0.9283
Motion Smoothness 0.9599 0.9691 0.9617
Dynamic Degree 0.4125 0.4088 0.4100
Overall Consistency 0.2529 0.2473 0.2509
5.3 Token Variations
@@ 6 Conclusion
Acknowledgement
We appreciate the fruitful discussions with Soobin Um. This work was supported by the National Research Foundation of Korea (NRF) under Grants RS-2024-00336454 and RS-2023-00262527, and by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program, KAIST).
@@ A.2 Hyperparameters for Evaluation
Tab. 5 lists the hyperparameters used for our quantitative evaluation. The ‘# of frames’ denotes the number of video frames decoded into pixels and used for optical flow calculation. Specifically, if the value is 6, this indicates that 3 sets of adjacent frames were sampled, resulting in 3 optical flows being used to compute the corresponding loss.
@@ A.3 User study
We conduct an A/B user study on Prolific, where 100 participants compared total 90 pairs of videos. Each participant was asked to choose the video they found more natural and smooth in motion by asking “Choose the more natural and smoothly moving video.”. The answer options were: A, B, or Both.
@@ B.1 Ablation on Token Optimization
Here, we conduct an ablation study to investigate the impact of various configurations related to token optimization on the overall performance (Tab. 6). Except for the factors being examined, all other settings were kept consistent with the values in Tab. 5.
@@
B.2 Synergies with the Existing Method
@@
While FreeInit significantly improves temporal quality, it does so at the expense of overall video quality. Specifically, Tab. 7 demonstrates that increasing the number of initialization steps leads to a sharp decline in metrics related to video generation quality, such as Overall Consistency and Image Quality. Fig. 6 further reveal that, compared to the baseline, the videos generated with FreeInit often lose high-frequency details and exhibit noticeable saturation. However, by applying our method after a single noise initialization step, we observed an improvement in temporal quality while relatively minimizing the compromise in video quality, compared to the FreeInit approach where noise initialization is applied multiple times.
B.3 Computational Cost
Model GPU Memory Computing Time
AnimateDiff 11.38 GiB 18.94 s/video
+Ours 35.60 GiB 50.94 s/video
Figure 8: Computational cost.
We compute GPU memory usage and computing time, averaging over 100 generations.